Beatrice invests $1,320 in an account that pays 4 percent simple interest. How much more could she have earned over a 5-year period if the interest had compounded annually?

A.) 32.00
B.) 21.98
C.) 25.60
D.) 18.50
Answer #1

Simple interest=1320*4%*5=264

compunded Interest=(1320*1.04^5)-1320=285.98

more interest=285.98-264=21.98